# Jarník-Besicovitch type theorems for semisimple algebraic groups

### v1

@@ 1.2. Main results
Let $\mathbf{P}_{0}$ be the minimal parabolic $\mathbb{Q}$-subgroup of $\mathbf{G}$ as defined above, and we may assume that the stable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in the unipotent radical $R_{u}(\mathbf{P}_{0})$ of $\mathbf{P}_{0}$ and its unstable horospherical subgroup is contained in the unipotent radical $R_{u}(\overline{\mathbf{P}}_{0})$ of $\overline{\mathbf{P}}_{0}$. The Bruhat decomposition of $\mathbf{G}$ is the following [6, §21]
$\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{W}}\cdot\mathbf{P}_{0}(\mathbb{Q})$
which implies that
$\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{W}}\cdot\overline{\mathbf{P}}_{0}(\mathbb{Q}),\quad\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{W}}\cdot R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{Q}).$
Let $\{w_{i}\}_{i\in I}$ be a set of representatives of ${}_{\mathbb{Q}}{W}$ in $\mathbf{G}(\mathbb{Q})$ where $I$ is the index set of ${}_{\mathbb{Q}}W$. For each $w_{i}$ $(i\in I)$, define
$\mathbf{F}_{w_{i}}=R_{u}(\overline{\mathbf{P}}_{0})\cap w_{i}^{-1}R_{u}(\overline{\mathbf{P}}_{\beta_{0}})w_{i},\quad\mathbf{H}_{w_{i}}=R_{u}(\overline{\mathbf{P}}_{0})\cap w_{i}^{-1}\mathbf{P}_{\beta_{0}}w_{i}$
and we have
$R_{u}(\overline{\mathbf{P}}_{0})=\mathbf{H}_{w_{i}}\cdot\mathbf{F}_{w_{i}}\quad(i\in I).$
Then one can further write the Bruhat decomposition as
$\mathbf{G}(\mathbb{Q})=$ $\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{W}}\cdot R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{Q})$
$=$ $\bigcup_{i\in I}\mathbf{P}_{0}(\mathbb{Q})\cdot(w_{i}\mathbf{H}_{w_{i}}w_{i}^{-1})(\mathbb{Q})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{Q})$
$=$ $\bigcup_{i\in I}\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{Q}).$
Note that the sets $\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{Q})$ $(i\in I)$ in the decomposition of $\mathbf{G}(\mathbb{Q})$ above may overlap, and for our purpose, we can choose any subset $\bar{I}$ of $I$ (as small as possible) such that
$\mathbf{G}(\mathbb{Q})=\bigcup_{i\in\bar{I}}\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{Q}).$
In the statement of Theorem 1.4 below, we will fix any such subset $\bar{I}$ in $I$.
Let $\bigwedge^{\dim V_{\beta_{0}}}V$ be the $\dim V_{\beta_{0}}$-exterior product vector space of $V$ over $\mathbb{C}$, and $\rho_{\beta_{0}}$ the natural extension of $\rho$ on $\bigwedge^{\dim V_{\beta_{0}}}V$. Let
@@
For each $i\in\bar{I}$, define the following morphism
$\Psi_{w_{i}}:w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1}(\mathbb{R})\to\bigwedge^{\dim V_{\beta_{0}}}V,\quad\Psi_{w_{i}}(x)=\rho_{\beta_{0}}(x)\cdot e_{V_{\beta_{0}}}.$
Note that $w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1}\subset R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ and $\Psi_{w_{i}}$ is an isomorphism onto its image. We denote by $a_{w_{i}}$ the growth rate of the asymptotic volume estimate of the real variety $\textup{Im}(\Psi_{w_{i}})$, and by $A_{w_{i}}$ the growth rate of the number of rational points in $w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1}(\mathbb{Q})$ (See §6 Corollary 6.8 and equation ($\ast$6) for more details). Let $\nu_{0}$ be the $\mathbb{Q}$-root in $R_{u}(\overline{\mathbf{P}}_{0})$ such that
@@
Now we can state the second main theorem in this note.
@@
Let $\mathbf{G}$ be a connected semisimple algebraic group defined over $\mathbb{Q}$, $\mathbf{T}$ a maximal $\mathbb{Q}$-split torus in $\mathbf{G}$ and $\{a_{t}\}_{t\in\mathbb{R}}$ a one-parameter subgroup in $\mathbf{T}(\mathbb{R})$. Let $\rho$ be a finite-dimensional irreducible representation of $\mathbf{G}$ defined over $\mathbb{Q}$ on a complex vector space $V$ with $\dim\ker\rho=0$. Then
$\dim_{H}S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})^{c}$
@@
$\leq$ $\max_{i\in\bar{I}}\left\{\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{w_{i}})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot\max\{a_{w_{i}},A_{w_{i}}\}\cdot\dim V_{\beta_{0}}\Big{|}\beta_{0}(w_{i}a_{-1}w_{i}^{-1})>\tau\right\}$
for any $0\leq\tau<\beta_{0}(a_{-1})$.
@@
It follows from the proof of Theorem 1.4 that for any $\tau\geq\beta_{0}(a_{-1})$, the subset $S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})^{c}=\emptyset$. (See §6.)
As corollaries of Theorems 1.3 and 1.4, we obtain the following results. The first result generalizes the Jarnik-Besicovitch type theorem on the homogneous space $X_{n}=\operatorname{\mathbf{SL}}_{n}(\mathbb{R})/\operatorname{\mathbf{SL}}_{n}(\mathbb{Z})$ for any one-parameter diagonal flow.

### v2

@@ 1.2. Main results
Let $\mathbf{P}_{0}$ be the minimal parabolic $\mathbb{Q}$-subgroup of $\mathbf{G}$ defined as in Theorem 1.4 such that the stable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in the unipotent radical $R_{u}(\mathbf{P}_{0})$ of $\mathbf{P}_{0}$ and the unstable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in the unipotent radical $R_{u}(\overline{\mathbf{P}}_{0})$ of $\overline{\mathbf{P}}_{0}$. It is known that the Bruhat decomposition of $\mathbf{G}$ is the following [8, Theorem 21.15]
$\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{W}}\cdot\mathbf{P}_{0}(\mathbb{Q}).$
Let $_{\mathbb{Q}}\mathcal{W}$} be a set of representatives of $_{\mathbb{Q}}{W}$ in $N(\mathbf{T})(\mathbb{Q})$. We may assume that $e\in{{}_{\mathbb{Q}}\mathcal{W}}$. Then one can deduce from the Bruhat decomposition that (see §2)
$\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{\mathcal{W}}}\cdot\overline{\mathbf{P}}_{0}(\mathbb{Q}),\quad\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{\mathcal{W}}}\cdot R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{Q}).$
For each $w\in{{}_{\mathbb{Q}}\mathcal{W}}$, define
$\mathbf{F}_{w}=R_{u}(\overline{\mathbf{P}}_{0})\cap w^{-1}R_{u}(\overline{\mathbf{P}}_{\beta_{0}})w,\quad\mathbf{H}_{w}=R_{u}(\overline{\mathbf{P}}_{0})\cap w^{-1}\mathbf{P}_{\beta_{0}}w$
and by [8, Proposition 21.9], we have
$R_{u}(\overline{\mathbf{P}}_{0})=\mathbf{H}_{w}\cdot\mathbf{F}_{w}.$
It follows that
$\mathbf{G}(\mathbb{Q})=$ $\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{\mathcal{W}}}\cdot R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{Q})$
$=$ $\bigcup_{w\in{{}_{\mathbb{Q}}\mathcal{W}}}\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot(w\mathbf{H}_{w}w^{-1})(\mathbb{Q})\cdot w\cdot\mathbf{F}_{w}(\mathbb{Q})$
$=$ $\bigcup_{w\in{{}_{\mathbb{Q}}\mathcal{W}}}\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w\cdot\mathbf{F}_{w}(\mathbb{Q}).$
Note that the subsets $\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w\cdot\mathbf{F}_{w}(\mathbb{Q})$ $(w\in{{}_{\mathbb{Q}}\mathcal{W}})$ above may overlap, and for our purpose, we choose a subset $_{\mathbb{Q}}\overline{\mathcal{W}}$} of $_{\mathbb{Q}}\mathcal{W}$} (as small as possible) such that
$\mathbf{G}(\mathbb{Q})=\bigcup_{w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}}\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w\cdot\mathbf{F}_{w}(\mathbb{Q}).$
In the statement of Theorem 1.5 below, we fix any such subset $_{\mathbb{Q}}\overline{\mathcal{W}}$} of $_{\mathbb{Q}}\mathcal{W}$}.
Let $\bigwedge^{\dim V_{\beta_{0}}}V$ be the $\dim V_{\beta_{0}}$-exterior product space of $V$ over $\mathbb{C}$, and denote by $\rho_{\beta_{0}}$ the natural extension of $\rho$ on $\bigwedge^{\dim V_{\beta_{0}}}V$. Let
@@
For each $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$, define the following morphism
$\Psi_{w}:\mathbf{F}_{w}(\mathbb{R})\to\bigwedge^{\dim V_{\beta_{0}}}V,\quad\Psi_{w}(x)=\rho_{\beta_{0}}(wxw^{-1})\cdot e_{V_{\beta_{0}}}.$
Note that $w\mathbf{F}_{w}w^{-1}\subset R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ and the stabilizer of the weight space $V_{\beta_{0}}$ is $\mathbf{P}_{\beta_{0}}$. So $\Psi_{w}$ is an isomorphism onto its image. For any $R>0$, let
$B_{R}=\left\{v\in\bigwedge^{\dim V_{\beta_{0}}}V:\|v\|\leq R\right\}$
where $\|\cdot\|$ is a Euclidean norm on $\bigwedge^{\dim V_{\beta_{0}}}V$. We denote by $a_{w}$ the growth rate of the asymptotic volume estimate of the real variety $\textup{Im}(\Psi_{w})$
$a_{w}=\lim_{R\to\infty}\frac{\log\mu_{\mathbf{F}_{w}}(\Psi_{w}^{-1}(B_{R}))}{\log R},$
where $\mu_{\mathbf{F}_{w}}$ is the Haar measure on $\mathbf{F}_{w}(\mathbb{R})$, and define by
$A_{w}=A_{w}(\Omega_{\beta_{0}}):=\limsup_{H\to\infty}\frac{\log|\{g\in\Omega_{\beta_{0}}\cap w\mathbf{F}_{w}w^{-1}(\mathbb{Q}):0pt(g)\leq H\}|}{\log H}$
the growth rate of the number of rational elements in $\Omega_{\beta_{0}}\cap w\mathbf{F}_{w}w^{-1}(\mathbb{Q})$ where $\Omega_{\beta_{0}}$ is a fixed open bounded subset in $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})$ containing a fundamental domain of $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})/(R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})\cap\Gamma)$ (See §6). The definition of the height function $0pt(\cdot)$ is given in §2 (Definition 2.5). As before, let $\nu_{0}$ be any $\mathbb{Q}$-root relative to $\mathbf{T}$ in $R_{u}(\overline{\mathbf{P}}_{0})$ such that
@@
Let $\mathbf{G}$ be a connected semisimple algebraic group defined over $\mathbb{Q}$, $\mathbf{T}$ a maximal $\mathbb{Q}$-split torus in $\mathbf{G}$ and $\{a_{t}\}_{t\in\mathbb{R}}$ a one-parameter subgroup in $\mathbf{T}(\mathbb{R})$. Let $\rho$ be a finite-dimensional irreducible representation of $\mathbf{G}$ defined over $\mathbb{Q}$ on a complex vector space $V$ with $\dim\ker\rho=0$, and $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ a function on $\mathbb{R}_{+}$. Then
$\dim_{H}S_{\rho}(\psi)^{c}$
@@
$\leq$ $\max_{\begin{subarray}{c}w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}\\ \beta_{0}(wa_{-1}w^{-1})\\ \geq\tau(\psi)\end{subarray}}\left\{\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(wa_{-1}w^{-1})-\tau(\psi))}{\nu_{0}(a_{1})}\cdot\max\{a_{w},A_{w}\}\cdot\dim V_{\beta_{0}}\right\}$
for any $0\leq\tau(\psi)\leq\beta_{0}(a_{-1})$.
Combining Theorem 1.4, Theorem 1.5 and [63, Theorem 4], we obtain the following theorem.
Theorem 1.6.
Let $\mathbf{G}$ be a connected $\mathbb{Q}$-simple group, $\mathbf{T}$ a maximal $\mathbb{Q}$-split torus in $\mathbf{G}$ and $\{a_{t}\}_{t\in\mathbb{R}}$ a one-parameter subgroup in $\mathbf{T}(\mathbb{R})$. Let $\rho$ be a finite-dimensional irreducible representation of $\mathbf{G}$ defined over $\mathbb{Q}$ on a complex vector space $V$ with $\dim\ker\rho=0$, and $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ a function on $\mathbb{R}_{+}$. Let $\beta_{0}$ and $\nu_{0}$ be defined as in Theorems 1.4 and 1.5. Suppose that the highest weight $\beta_{0}$ is a multiple of $\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}{\alpha}$. Then
$\dim_{H}S_{\rho}(\psi)^{c}=$ $\dim\mathbf{G}-\frac{\tau(\psi)}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}\cdot\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}\alpha(a_{1})$
for any $0\leq\tau(\psi)<\beta_{0}(a_{-1})$. In the case $\tau(\psi)=\beta_{0}(a_{-1})$, we have
  1. (1)
    If $\psi(t)\cdot e^{\beta_{0}(a_{-1})t}$ is bounded, then $S_{\rho}(\psi)^{c}=\emptyset$.
  2. (2)
    If $\psi(t)\cdot e^{\beta_{0}(a_{-1})t}$ is unbounded, then $S_{\rho}(\psi)^{c}\neq\emptyset$ and
    $\dim_{H}S_{\rho}(\psi)^{c}=\dim\mathbf{G}-\frac{\tau(\psi)}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}\cdot\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}\alpha(a_{1}).$
Any irreducible $\mathbb{Q}$-rational representation of a $\mathbb{Q}$-rank one $\mathbb{Q}$-simple group satisfies the conditions of Theorem 1.6. In this case, Theorem 1.6 recovers and largely extends [82, Theorem 1.1] in the arithmetic setting. We will discuss the relation between Theorem 1.6 and [82, Theorem 1.1] in §7. We remark that in [82], we also deal with the non-arithmetic case due to the fact that there are non-arithmetic lattices in $\mathbb{R}$-rank one simple groups.
Besides $\mathbb{Q}$-rank one $\mathbb{Q}$-simple groups, in the following we also list some other examples of Theorem 1.6. The first example concerns the standard representation of $\operatorname{\mathbf{SL}}_{n}$ and any one-parameter diagonal subgroup $\{a_{t}\}_{t\in\mathbb{R}}$ in $\operatorname{\mathbf{SL}}_{n}(\mathbb{R})$. It can be considered as a generalization of [11, 20] from the viewpoint of shrinking target problem.
Theorem 1.7.
Let $\rho:\operatorname{\mathbf{SL}}_{n}\to\operatorname{\mathbf{GL}}(V)$ be the standard representation of $\operatorname{\mathbf{SL}}_{n}$ on $V=\mathbb{C}^{n}$ defined via matrix multiplication
$\rho(g)\cdot v=g\cdot v\quad(g\in\operatorname{\mathbf{SL}}_{n},v\in V).$
Let $\{a_{t}\}_{t\in\mathbb{R}}$ be a one-parameter diagonal subgroup in $\operatorname{\mathbf{SL}}_{n}(\mathbb{R})$, $\beta_{0}$ the highest weight of $\rho$ with respect to $\{a_{t}\}_{t\in\mathbb{R}}$ and $\nu_{0}$ the $\mathbb{Q}$-root in $\operatorname{\mathbf{SL}}_{n}$ defined as in Theorem 1.6. Let $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ be a function on $\mathbb{R}_{+}$. Then we have
$\dim_{H}S_{\rho}(\psi)^{c}=\dim\operatorname{\mathbf{SL}}_{n}-\frac{n\cdot\tau(\psi)}{\nu_{0}(a_{1})}$
for any $0\leq\tau(\psi)<\beta_{0}(a_{-1})$. In the case $\tau(\psi)=\beta_{0}(a_{-1})$, we have
  1. (1)
    If $\psi(t)\cdot e^{\beta_{0}(a_{-1})t}$ is bounded, then $S_{\rho}(\psi)^{c}=\emptyset$.
  2. (2)
    If $\psi(t)\cdot e^{\beta_{0}(a_{-1})t}$ is unbounded, then $S_{\rho}(\psi)^{c}\neq\emptyset$ and
    $\dim_{H}S_{\rho}(\psi)^{c}=\dim\operatorname{\mathbf{SL}}_{n}-\frac{n\cdot\tau(\psi)}{\nu_{0}(a_{1})}.$
The second example deals with the adjoint representation of $\operatorname{\mathbf{SL}}_{n}$. We will prove in §7 that it generalizes the main result of [25].
Theorem 1.8.
Let $\rho:\operatorname{\mathbf{SL}}_{n}\to\operatorname{\mathbf{GL}}(V)$ be the adjoint representation of $\operatorname{\mathbf{SL}}_{n}$ where $V=\mathfrak{sl}_{n}$ is the Lie algebra of $\operatorname{\mathbf{SL}}_{n}$. Let $\{a_{t}\}_{t\in\mathbb{R}}$ be a one-parameter diagonal subgroup in $\operatorname{\mathbf{SL}}_{n}(\mathbb{R})$, $\nu_{0}$ the $\mathbb{Q}$-root in $\operatorname{\mathbf{SL}}_{n}$ defined as in Theorem 1.6 and $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ a function on $\mathbb{R}_{+}$. Then we have
$\dim_{H}S_{\rho}(\psi)^{c}=\dim\operatorname{\mathbf{SL}}_{n}-\frac{(n-1)\cdot\tau(\psi)}{\nu_{0}(a_{1})}$
for any $0\leq\tau(\psi)<\nu_{0}(a_{1})$. In the case $\tau(\psi)=\nu_{0}(a_{1})$, we have
  1. (1)
    If $\psi(t)\cdot e^{\nu_{0}(a_{1})t}$ is bounded, then $S_{\rho}(\psi)^{c}=\emptyset$.
  2. (2)
    If $\psi(t)\cdot e^{\nu_{0}(a_{1})t}$ is unbounded, then $S_{\rho}(\psi)^{c}\neq\emptyset$ and
    $\dim_{H}S_{\rho}(\psi)^{c}=\dim\operatorname{\mathbf{SL}}_{n}-\frac{(n-1)\cdot\tau(\psi)}{\nu_{0}(a_{1})}.$
@@
One can see from the proofs that all the theorems stated above also hold for $\dim_{H}(S_{\rho}(\psi)^{c}\cap U)$ where $U$ is any open bounded subset in $\mathbf{G}(\mathbb{R})$. (See Remark 5.5.)
1.3. Connections to Diophantine approximation
In this subsection, we discuss two connections of Theorem 1.6 to the metric theory of Diophantine approximation on certain algebraic varieties, and present some related Jarník-Besicovitch type theorems.
1.3.1. Connection to Diophantine approximation on flag varieties
Diophantine approximation on group varieties is an open problem raised by Lang [56]. In recent years, many important results have been established about Diophantine approximation on various group varieties including spheres, quadrics and algebraic groups [28, 29, 22, 31, 1, 18, 51, 21, 52, 30, 27]. Here we discuss a connection between Theorem 1.6 and Diophantine approximation on generalized flag varieties, and derive an analogue of Jarník-Besicovitch theorem. We mainly follow the exposition in [17] by de Saxcé. For more details about Diophantine approximation on flag varieties, one may refer to [17].
Let $\mathbf{X}=\mathbf{P}\backslash\mathbf{G}$ be a flag variety where $\mathbf{G}$ is a connected $\mathbb{Q}$-simple group and $\mathbf{P}$ a parabolic $\mathbb{Q}$-subgroup of $\mathbf{G}$. Let $V$ be a finite dimensional vector space defined over $\mathbb{Q}$, $\|\cdot\|$ a Euclidean norm on $V$ and $(e_{i})_{1\leq i\leq d}$ a rational orthonormal basis of $V$. For any rational point $v$ in $\mathbf{P}(V)(\mathbb{Q})$, we define the height $H(v)$ of $v$ by
$H(v)=\|\mathbf{v}\|$
where $\mathbf{v}$ is the representative primitive integral vector (up to sign) of $v$ in $\oplus_{1\leq i\leq d}\mathbb{Z}e_{i}$. Note that the height function $H(\cdot)$ is defined up to a multiplicative constant.
If $\rho_{\chi}:\mathbf{G}\to\operatorname{\mathbf{GL}}(V)$ is an irreducible $\mathbb{Q}$-rational representation of $\mathbf{G}$ generated by a one-dimensional weight space $V_{\chi}\subset V$ of highest weight $\chi$ such that the stabilizer of $V_{\chi}$ is $\mathbf{P}$, then we may identify $\mathbf{X}=\mathbf{P}\backslash\mathbf{G}$ with orbit $\mathbf{G}\cdot V_{\chi}$ in $\mathbf{P}(V)$ by
$\mathbf{X}\to\mathbf{G}\cdot V_{\chi},\quad\mathbf{P}\cdot g\mapsto\rho_{\chi}(g^{-1})\cdot V_{\chi}\in\mathbf{G}\cdot V_{\chi}$
and the height of any rational point in $\mathbf{X}=\mathbf{P}\backslash\mathbf{G}$ can be defined by restriction of the height function $H(\cdot)$ on $\mathbf{G}\cdot V_{\chi}\subset\mathbf{P}(V)$.
As in Theorem 1.6, we may choose a maximal $\mathbb{Q}$-split torus $\mathbf{T}$ and a minimal parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{0}$ in $\mathbf{G}$ such that $\mathbf{T}\subset\mathbf{P}_{0}\subset\mathbf{P}$. Then $\mathbf{P}_{0}$ and $\mathbf{T}$ define a root system ($\Phi,\;\Phi^{+},\;\Delta$) where $\Phi$ is the set of $\mathbb{Q}$-roots relative to $\mathbf{T}$, $\Phi^{+}$ is the set of positive $\mathbb{Q}$-roots determined by $\mathbf{P}_{0}$ and $\Delta$ is the set of simple $\mathbb{Q}$-roots in $\Phi^{+}$. We denote by $R_{u}(\mathbf{P})$ (resp. $R_{u}(\mathbf{P}_{0})$) the unipotent radical of $\mathbf{P}$ (resp. $\mathbf{P}_{0}$), and write $\Phi(R_{u}(\mathbf{P}))$ for the set of $\mathbb{Q}$-roots in $R_{u}(\mathbf{P})$ relative to $\mathbf{T}$. The symbol $\sum_{\alpha\in\Phi(R_{u}(\mathbf{P}))}$ stands for the sum over all $\mathbb{Q}$-roots $\alpha\in\Phi(R_{u}(\mathbf{P}))$ counted with multiplicities (i.e., the dimensions of the $\mathbb{Q}$-root spaces associated to $\alpha\in\Phi(R_{u}(\mathbf{P}))$ in the Lie algebra of $R_{u}(\mathbf{P})$).
In order to apply Theorem 1.6, we choose a one-parameter subgroup $\{a_{t}\}_{t\in\mathbb{R}}$ in $\mathbf{T}(\mathbb{R})$ as follows: $a_{t}=\exp(t\cdot Y)$ where $Y$ is an element in the Lie algebra of $\mathbf{T}(\mathbb{R})$ satisfying
$\begin{cases}\alpha(Y)=\alpha(a_{1})=0,\textup{ if }\alpha\in\Delta_{\mathbf{P}}\\ \alpha(Y)=\alpha(a_{1})=-1,\textup{ if }\alpha\in\Delta\setminus\Delta_{\mathbf{P}}\end{cases}.$
Here $\Delta_{\mathbf{P}}\subset\Delta$ is the set of simple $\mathbb{Q}$-roots associated to $\mathbf{P}$ . We denote by $\nu_{0}$ any $\mathbb{Q}$-root relative to $\mathbf{T}$ in $\Phi^{+}$ satisfying
$\nu_{0}(a_{-1})=\max\{\alpha(a_{-1}):\alpha\in\Phi^{+}\}.$
On the variety $\mathbf{X}(\mathbb{R})$, there is a Carnot-Carathéodory distance $d_{\textup{CC}}$, and its explicit definition can be found in [17, Chapter 2, §2]. With this Carnot-Carathéodory distance, for any $x\in\mathbf{X}(\mathbb{R})$, one may define the Diophantine exponent $\beta_{\chi}(x)$ of $x$ by
$\beta_{\chi}(x)=\inf\{\beta>0:\exists\textup{$C>0$ such that }d_{\textup{CC}}(x,v)\geq C\cdot H(v)^{-\beta}\;(\forall v\in\mathbf{X}(\mathbb{Q}))\}.$
We have the following theorem about the Diophantine exponent $\beta_{\chi}(x)$.
Theorem 1.10 ([17, Theorem 2.4.5]).
There exits a positive constant $\beta_{\chi}>0$ such that for almost every $x\in\mathbf{X}(\mathbb{R})$ (with respect to a Riemannian volume form on $\mathbf{X}(\mathbb{R})$)
$\beta_{\chi}(x)=\beta_{\chi}.$
Moreover, we have $\beta_{\chi}=-1/\chi(Y)=1/\chi(a_{-1})$.
Definition 1.11.
For any decreasing function $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$, define the set of $\psi$-Diophantine points in $\mathbf{X}(\mathbb{R})$ by
$E_{\mathbf{X}}(\psi)=\{x\in\mathbf{X}(\mathbb{R}):\exists C>0\textup{ such that }d_{\textup{CC}}(x,v)\geq C\cdot(H(v))^{-\beta_{\chi}}\psi(H(v))\;(\forall v\in\mathbf{X}(\mathbb{Q}))\}$
where the constant $\beta_{\chi}$ is defined as in Theorem 1.10. We denote by $E_{\mathbf{X}}(\psi)^{c}$ the complement of $E_{\mathbf{X}}(\psi)$ in $\mathbf{X}(\mathbb{R})$.
For any function $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$, define the lower order at infinity of $\psi$ by
$\gamma(\psi)=\liminf_{t\to\infty}\left(-\frac{\ln\psi(t)}{\ln t}\right).$
From a correspondence theorem [17, Proposition 3.2.4] and Theorem 1.6, we may deduce the following result.